# Diode effect in the Fraunhofer pattern of disordered planar Josephson junctions

### v2

@@ II.1 Absence of JDE in mirror-symmetric junctions
It is crucial to appreciate how other choices of zero-field mirror symmetry cannot yield conclusive statements about the absence of JDE. For example, for a system whose scattering matrix satisfies $M_{y}s(\Phi)M_{y}^{-1}=s(-\Phi)$, we obtain $M_{y}A(\varphi,\Phi)M_{y}^{-1}=r_{A}^{*}(\varphi,\Phi)r_{A}^{*}$, so that we cannot conclude anything regarding the presence or absence of JDE. Furthermore, it is interesting to consider the inversion symmetry, or parity, $\hat{P}$, which maps ${\bf r}\to-{\bf r}$. Inversion symmetry can be decomposed as the product of three orthogonal mirror symmetries, $\hat{P}=M_{x}M_{y}M_{z}$. Assuming that $M_{y}$ is broken implies breaking of the inversion symmetry. At the same time, a zero-field mirror symmetry about $M_{x}$ rules out the JDE, showing that, generally, inversion symmetry breaking is only a necessary condition.
@@ II.2 Planar Josephson junctions
where $c_{i\sigma}$ describes electrons with spin $\sigma=\uparrow,\downarrow$ at position ${\bf r}_{i}$, and the index $i$ runs over all the lattice points of the central region. The square lattice has a sublattice symmetry that yields a particle-hole symmetry in the spectrum, so that positive and negative values of the chemical potential yield the same results. Electrons have onsite energy $U_{i}=4t-\mu+\delta U_{i}$, comprising a chemical potential contribution $\mu$ and an additional onsite potential $\delta U_{i}\equiv\delta U({\bf r}_{i})$, and can hop among nearest-neighboring sites with amplitude $te^{i\theta_{i,j}}$, where the Peierls phase $\theta_{i,j}=\frac{e}{\hbar}\int_{{\bf r}_{i}}^{{\bf r}_{j}}{\bf A}({\bf r})\cdot d{\bf s}$ is due to the external magnetic field ${\bf B}=\nabla\times{\bf A}$. The central region is attached to the superconducting leads, and potential barriers that shift local energy are added to the contact interface. We consider transport at zero energy in the limit in which the energy dependence of the scattering matrix can be neglected, that corresponds to a system length $L<\xi$, with $\xi$ the coherence length of the superconducting leads, and numerically simulate several systems employing the KWANT package [42] (see  [^2] Supplementary Material: Sec. II - Comparison with realistic systems).
@@ II.2.1 Smooth Potentials
Figure 4: Fraunhofer interference pattern of a planar Josephson junction with short-range disorder. The junction structure is as in Fig. 2. (a) Critical currents $I_{c}^{\pm}$ as a function of the external flux $\Phi$ resulting from the local onsite potential from the local onsite potential $\delta U_{i}$, with $U_{0}/2<\delta U_{i}<U_{0}/2$ randomly distributed for $U_{0}=0.1~t$, shown in the top-right inset. (b) Rectification coefficient $\eta$ for the critical currents shown in (a). (c) Critical currents $I_{c}^{\pm}$ as a function of the external flux $\Phi$ resulting from a local onsite potential $\delta U_{i}$, with $U_{0}/2<\delta U_{i}<U_{0}/2$ randomly distributed for $U_{0}=0.1~t$ and symmetrized $\delta U_{i}\to(\delta U_{i}+M_{x}\delta U_{i}M_{x}^{-1})/2$, shown in the top-right inset. (d) Rectification coefficient $\eta$ for the critical currents shown in (c). The chemical potential is set to $\mu=0.3~t$.
@@
We first consider a central rectangular region of width W=100aW=100~a and length L=20aL=20~a, in terms of a microscopic discretization length $a$, schematized in Fig. 2(f). We set the chemical potential at $\mu=0.3~t$ in both the central region and in the leads and apply a non-uniform potential $\delta U(x,y)=V_{g}(\tanh((x-W/2)/L_{s})+1)/2$. The potential, which also comprises a shift of the chemical potential, is not symmetric under $M_{x}$, such that $M_{x}\delta U(x,y)M_{x}^{-1}\neq\delta U(-x,y)$, and it is shown in the upper right inset of Fig. 2(a) for a value of $V_{g}=0.1~t$, producing a smooth profile that changes in length $L_{s}=20~a$. Potential barriers that shift the local energy by 0.2t0.2~t are added at the interface with the contacts. In Fig. 2(a), we show the associated Fraunhofer interference pattern for the critical currents $I_{c}^{\pm}$. We see that the presence of a smooth potential breaking the mirror symmetry $M_{x}$ produces a diode effect, as quantified by the rectification coefficient $\eta$ shown in Fig. 2(b), reaching rectification $30\%$.
@@
The fact that a smooth, long-wavelength potential can yield a diode effect suggests that the latter can be tuned by applying a gate. The potential shown in the inset of Fig. 2(a) well approximates the one produced by a back gate that occupies the right side of the junction. In Fig. 2 (c), we plot the rectification efficiency for a few values of the gate potential, $V_{g}/t=0.02,0.04,0.06,0.08$. In the first lobe we clearly see a trend for which the rectification increases with the gate potential. In turn, we see no clear systematic evolution of the curves for increasing gate strength. The rectifications shows spikes in correspondence of the nodes of the Fraunhofer pattern, which are due destructive interference acting in different ways for trajectories that go from one lead to the other and vice versa.
@@
We then consider a potential that is even under $M_{x}$ but that is not symmetric under $M_{y}$, thus breaking the inversion symmetry $\hat{P}$. Specifically, we choose $\delta U(x,y)=V\tanh((y-L/2)/L_{s})$. The potential satisfies $M_{x}\delta U(x,y)M_{x}^{-1}=\delta U(-x,y)$, $M_{y}\delta U(x,y)M_{y}^{-1}\neq\delta U(x,-y)$, and it is shown in the top-right inset of Fig. 2(d) for a value of $V_{g}=0.1~t$, yielding a smooth profile that changes on a length $L_{s}=20~a$. As clearly shown in Fig. 2(d), the two curves $I^{\pm}_{c}(\Phi)$ cannot be distinguished and the rectification coefficients are less than $0.4\%$, as shown in Fig. 2(e). This result is significant, as it corroborates the finding that breaking of inversion symmetry cannot guarantee an JDE and that the presence of a symmetric profile under $M_{x}$ guarantees no JDE.
@@ II.2.3 Short wavelength disorder
We then consider the case iii) of a potential that varies on a short-length scale, with local impurities of strength $\delta U_{i}$ randomly distributed between $U_{0}/2<\delta U_{i}<U_{0}/2$, as shown in the upper left panel of Fig. 4 for $U_{0}=0.1~t$. This is a relatively weak short-wavelength disorder of the Anderson type, that yields a long mean-free path $\ell_{\rm mf}$, much larger than the Fermi wavelength. The resulting Fraunhofer interference pattern for the critical currents $I_{c}^{\pm}$ is shown in Fig. 4(a), together with the rectification coefficient in Fig. 4(b). Analogously to the previous case, the broken mirror symmetry $M_{x}$ is sufficient to give a diode effect in the Fraunhofer interference pattern, which is qualitatively very similar to the case of a smooth uniform potential of Fig. 2(a,b). In particular, spikes appear in the rectification in correspondence of the nodes of the Fraunhofer pattern that are accompanied by sign changes of the rectification. In addition, we check that by imposing symmetry under $M_{x}$ the JDE disappears. To this end, we first generate a configuration with $\delta U_{i}$ randomly distributed between $U_{0}/2<\delta U_{i}<U_{0}/2$, and then symmetrize it such that $\delta U_{i}\to(\delta U_{i}+M_{x}\delta U_{i}M_{x}^{-1})/2$. The resulting potential is shown in the upper right panel of Fig. 4(c) and the associated currents $I_{c}^{\pm}$ shown in Fig. 3(c) are not distinguishable with the naked eye, as can be checked in Fig. 4(d) where a maximal rectification of $0.4\%$ appears.
@@ II.3 Disorder and Fluctuations
Figure 5: Statistical analysis of disordered configurations. (a) Rectification coefficient $\eta$ for 50 disorder configurations and a given value of the strength $U_{0}=1.2~t$. In black is the average rectification of the 50 random disorder configurations. (b) Root mean square of the rectification coefficient, ${\rm rms}~\eta\equiv\sqrt{\langle\eta^{2}\rangle-\langle\eta\rangle^{2}}$, versus the applied flux, averaged over 50 different pseudorandom disorder configurations and different values of $U_{0}$. The curve corresponding to $U_{0}=1.2~t$ is the square root of the average of the square of the curves in (a). In (a) and (b) $\mu=0.3~t$, $L=20$, $W=100$ and there is no barrier. (c) ${\rm rms}~\eta\times(t/U_{0})$ providing a scaling of the curves in (b) with the impurity strength.
@@
These signatures are clear manifestations of the fluctuations of the critical current due to disorder and magnetic field. At the same time, we recall that at finite magnetic field, the current cannot be expressed as a linear statistics in the transmission eigenvalues. A common trend emerges in Fig. 5(c), which requires further study. At the same time, a universal behavior emerges only qualitatively, because the strength of the disorder quantifies the mirror symmetry breaking character of the system and thus the resulting JDE properties. We rather notice that the spikes at the node of the Fraunhofer pattern are a clear manifestation of the mirror symmetry breaking and can help in distinguishing other intrinsic sources of JDE.
@@ II.4 Multiple-loop SQUID
where $\epsilon_{n,j}$ are the Foruier component of the EPR of the $j$-th junction [49], $I_{0}=2\pi\Delta/\Phi_{0}$ and $\varphi_{j}=2\pi BA_{j}/\Phi_{0}$ is the flux through the area $A_{j}$ of the loop between junctions j1j-1 and $j$, in units of $\Phi_{0}/2\pi$. The expression Eq. (11) meets two fundamental conditions at the same time: i) it breaks time-reversal symmetry, and ii) it generally breaks mirror symmetry $j\to N+1-j$, thus allowing, in general, the appearance of a diode effect. In addition, the transmissions $\tau_{j}$ need to be sizable. This can be accomplished in two ways, i) by considering different transparencies $\tau_{j}$ and same areas of the loops, for which $\varphi_{j}=2\pi(j-1)\Phi/N\Phi_{0}$, or ii) by keeping the same transparency $\tau$ for the junctions and considering different areas $A_{j}$ of the loops, such that $\varphi_{j}=2\pi\sum_{i\leq j}A_{j}B/\Phi_{0}$.
@@ III Discussion
The presented analysis sheds light on the ubiquity of the diode effect in wide planar Josephson junctions, showing how seemingly different spatially nonuniform junctions yield qualitatively similar JDE. The spatial dependence of the phase that Cooper pairs acquire in an orbital magnetic field fully emerges in the Fraunhofer interference pattern in planar Josephson junctions, which turns out to be particularly suitable to highlight the relevant symmetries enabling the onset of JDE. In particular, the conditions of time-reversal and inversion symmetry breaking are known to be necessary conditions that do not guarantee the onset of a diode effect. In turn, casting the analysis in terms of mirror symmetry breaking is more stringent. Although it is only possible to demonstrate that junctions that do not break the relevant mirror symmetry cannot show JDE, all the cases studied do indeed show JDE.
@@ IV Acknowledgments
We acknowledge S. Heun for valuable discussions and for carefully reading the manuscript. L.C. acknowledges the Fondazione Cariplo under the grant 2023-12594. M.C. and F.G. acknowledge the EU’s Horizon 2020 Research and Innovation Framework Programme under Grants No. 964398 (SUPERGATE) and the PNRR MUR project PE0000023-NQSTI for partial financial support. F.G. acknowledges partial financial support under Grant No. 101057977 (SPECTRUM). E.S. acknowledges the project HelicS, project number DFM.AD002.206. We acknowledge NextGenerationEU PRIN project 2022A8CJP3 (GAMESQUAD) for partial financial support.

### v3

@@ II.1 Absence of JDE in mirror-symmetric junctions
It is crucial to appreciate how other choices of zero-field mirror symmetry cannot yield conclusive statements about the absence of JDE. For example, for a system whose scattering matrix satisfies $M_{y}s(\Phi)M_{y}^{-1}=s(-\Phi)$, we obtain $M_{y}A(\varphi,\Phi)M_{y}^{-1}=r_{A}^{*}A(\varphi,\Phi)r_{A}^{*}$, so that we cannot conclude anything regarding the presence or absence of JDE. Furthermore, it is interesting to consider the inversion symmetry, or parity, $\hat{P}$, which maps ${\bf r}\to-{\bf r}$. Inversion symmetry can be decomposed as the product of three orthogonal mirror symmetries, $\hat{P}=M_{x}M_{y}M_{z}$. Assuming that $M_{y}$ is broken implies breaking of the inversion symmetry. At the same time, a zero-field mirror symmetry about $M_{x}$ rules out the JDE, showing that, generally, inversion symmetry breaking is only a necessary condition.
@@ II.2 Planar Josephson junctions
where $c_{i\sigma}$ describes electrons with spin $\sigma=\uparrow,\downarrow$ at position ${\bf r}_{i}$, and the index $i$ runs over all the lattice points of the central region. The square lattice has a sublattice symmetry that yields a particle-hole symmetry in the spectrum. Electrons have onsite energy $U_{i}=4t-\mu+\delta U_{i}$, comprising a chemical potential contribution $\mu$ and an additional onsite potential $\delta U_{i}\equiv\delta U({\bf r}_{i})$, and can hop among nearest-neighboring sites with amplitude $te^{i\theta_{i,j}}$, where the Peierls phase $\theta_{i,j}=\frac{e}{\hbar}\int_{{\bf r}_{i}}^{{\bf r}_{j}}{\bf A}({\bf r})\cdot d{\bf s}$ is due to the external magnetic field ${\bf B}=\nabla\times{\bf A}$. The central region is attached to the superconducting leads, and potential barriers that shift local energy are added to the contact interface. We consider transport at zero energy in the limit in which the energy dependence of the scattering matrix can be neglected, that corresponds to a system length $L<\xi$, with $\xi$ the coherence length of the superconducting leads, and numerically simulate several systems employing the KWANT package [42] (see  [^2] Supplementary Material: Sec. II - Comparison with realistic systems).
@@ II.2.1 Smooth Potentials
Figure 4: Fraunhofer interference pattern of a planar Josephson junction with short-range disorder. The junction structure is as in Fig. 2. (a) Critical currents $I_{c}^{\pm}$ as a function of the external flux $\Phi$ resulting from the local onsite potential $\delta U_{i}$, with $U_{0}/2<\delta U_{i}<U_{0}/2$ randomly distributed for $U_{0}=0.1~t$, shown in the top-right inset. (b) Rectification coefficient $\eta$ for the critical currents shown in (a). (c) Critical currents $I_{c}^{\pm}$ as a function of the external flux $\Phi$ resulting from a local onsite potential $\delta U_{i}$, with $U_{0}/2<\delta U_{i}<U_{0}/2$ randomly distributed for $U_{0}=0.1~t$ and symmetrized $\delta U_{i}\to(\delta U_{i}+M_{x}\delta U_{i}M_{x}^{-1})/2$, shown in the top-right inset. (d) Rectification coefficient $\eta$ for the critical currents shown in (c). The chemical potential is set to $\mu=0.3~t$.
@@
We first consider a central rectangular region of width W=100aW=100~a and length L=20aL=20~a, in terms of a microscopic discretization length $a$, schematized in Fig. 2(f). We set the chemical potential at $\mu=0.3~t$ in both the central region and in the leads and apply a non-uniform potential $\delta U(x,y)=V_{g}(\tanh((x-W/2)/L_{s})+1)/2$. The potential, which also comprises a shift of the chemical potential, is not symmetric under $M_{x}$, such that $M_{x}\delta U(x,y)M_{x}^{-1}\neq\delta U(-x,y)$, and it is shown in the upper right inset of Fig. 2(a) for a value of $V_{g}=0.1~t$, producing a smooth profile that changes on a length $L_{s}=20~a$. Potential barriers that shift the local energy by 0.2t0.2~t are added at the interface with the contacts. In Fig. 2(a), we show the associated Fraunhofer interference pattern for the critical currents $I_{c}^{\pm}$. We see that the presence of a smooth potential breaking the mirror symmetry $M_{x}$ produces a diode effect, as quantified by the rectification coefficient $\eta$ shown in Fig. 2(b), reaching rectification $30\%$.
@@
The fact that a smooth, long-wavelength potential can yield a diode effect suggests that the latter can be tuned by applying a gate. The potential shown in the inset of Fig. 2(a) well approximates the one produced by a back gate that occupies the right side of the junction. In Fig. 2 (c), we plot the rectification efficiency for few values of the gate potential, $V_{g}/t=0.02,0.04,0.06,0.08$. In the first lobe we clearly see a trend for which the rectification increases with the gate potential. In turn, we see no clear systematic evolution of the curves for increasing gate strength. The rectifications shows spikes in correspondence of the nodes of the Fraunhofer pattern, which are due destructive interference acting in different ways for trajectories that go from one lead to the other and vice versa.
@@
We then consider a potential that is even under $M_{x}$ but that is not symmetric under $M_{y}$, thus breaking the inversion symmetry $\hat{P}$. Specifically, we choose $\delta U(x,y)=V\tanh((y-L/2)/L_{s})$. The potential satisfies $M_{x}\delta U(x,y)M_{x}^{-1}=\delta U(-x,y)$, $M_{y}\delta U(x,y)M_{y}^{-1}\neq\delta U(x,-y)$, and it is shown in the top-right inset of Fig. 2(d) for a value of $V_{g}=0.1~t$, yielding a smooth profile that changes on a length $L_{s}=20~a$. As clearly shown in Fig. 2(d), the two curves $I^{\pm}_{c}(\Phi)$ cannot be distinguished and the rectification coefficients are less than $0.4\%$, as shown in Fig. 2(e). This result is significant, as it corroborates the finding that breaking of inversion symmetry cannot guarantee a JDE and that the presence of a symmetric profile under $M_{x}$ guarantees no JDE.
@@ II.2.3 Short wavelength disorder
We then consider the case iii) of a potential that varies on a short length scale, with local impurities of strength $\delta U_{i}$ randomly distributed between $U_{0}/2<\delta U_{i}<U_{0}/2$, as shown in the upper left panel of Fig. 4 for $U_{0}=0.1~t$. This is a relatively weak short-wavelength disorder of the Anderson type, that yields a long mean-free path $\ell_{\rm mf}$, much larger than the Fermi wavelength. The resulting Fraunhofer interference pattern for the critical currents $I_{c}^{\pm}$ is shown in Fig. 4(a), together with the rectification coefficient in Fig. 4(b). Analogously to the previous case, the broken mirror symmetry $M_{x}$ is sufficient to give a diode effect in the Fraunhofer interference pattern, which is qualitatively very similar to the case of a smooth uniform potential of Fig. 2(a,b). In particular, spikes appear in the rectification in correspondence of the nodes of the Fraunhofer pattern that are accompanied by sign changes of the rectification. In addition, we check that by imposing symmetry under $M_{x}$ the JDE disappears. To this end, we first generate a configuration with $\delta U_{i}$ randomly distributed between $U_{0}/2<\delta U_{i}<U_{0}/2$, and then symmetrize it such that $\delta U_{i}\to(\delta U_{i}+M_{x}\delta U_{i}M_{x}^{-1})/2$. The resulting potential is shown in the upper right panel of Fig. 4(c) and the associated currents $I_{c}^{\pm}$ shown in Fig. 3(c) are not distinguishable with the naked eye, as can be checked in Fig. 4(d) where a maximal rectification of $0.4\%$ appears.
@@ II.3 Disorder and Fluctuations
Figure 5: Statistical analysis of disordered configurations. (a) Rectification coefficient $\eta$ for 50 disorder configurations and a given value of the strength $U_{0}=1.2~t$. In black is the average rectification of the 50 random disorder configurations. (b) Root mean square of the rectification coefficient, ${\rm rms}~\eta\equiv\sqrt{\langle\eta^{2}\rangle-\langle\eta\rangle^{2}}$, averaged over 50 different pseudorandom disorder configurations, versus the applied flux for different values of $U_{0}$. The curve corresponding to $U_{0}=1.2~t$ is the square root of the average of the square of the curves in (a). In (a) and (b) $\mu=0.3~t$, $L=20$, $W=100$ and there is no barrier. (c) ${\rm rms}~\eta\times(t/U_{0})$ providing a scaling of the curves in (b) with the impurity strength.
@@
These signatures are clear manifestations of the fluctuations of the critical current due to disorder and magnetic field. At the same time, we recall that at finite magnetic field, the current cannot be expressed as a linear statistics in the transmission eigenvalues. A common trend emerges in Fig. 5(c), which points to a qualitatively universal behaviour that requires further study. We further notice that the spikes at the node of the Fraunhofer pattern are a clear manifestation of the mirror symmetry breaking and can help in distinguishing other intrinsic sources of JDE.
@@ II.4 Multiple-loop SQUID
where $\epsilon_{n,j}$ are the Foruier component of the EPR of the $j$-th junction [49], $I_{0}=2\pi\Delta/\Phi_{0}$ and $\varphi_{j}=2\pi BA_{j}/\Phi_{0}$ is the flux through the area $A_{j}$ of the loop between junctions j1j-1 and $j$, in units of $\Phi_{0}/2\pi$. The expression Eq. (11) meets two fundamental conditions at the same time: i) it breaks time-reversal symmetry, and ii) it generally breaks mirror symmetry $j\to N+1-j$, thus allowing, in general, the appearance of a diode effect. This can be accomplished in two ways, i) by considering different transparencies $\tau_{j}$ and same areas of the loops, for which $\varphi_{j}=2\pi(j-1)\Phi/N\Phi_{0}$, or ii) by keeping the same transparency $\tau$ for the junctions and considering different areas $A_{j}$ of the loops, such that $\varphi_{j}=2\pi\sum_{i\leq j}A_{j}B/\Phi_{0}$. In addition, the transmissions $\tau_{j}$ need to be sizable.
@@ III Discussion
The presented analysis sheds light on the ubiquity of the diode effect in wide planar Josephson junctions, showing how seemingly different spatially nonuniform junctions yield qualitatively similar JDE. The spatial dependence of the phase that Cooper pairs acquire in an orbital magnetic field fully emerges in the Fraunhofer interference pattern in planar Josephson junctions, which turn out to be particularly suitable to highlight the relevant symmetries enabling the onset of JDE. In particular, the conditions of time-reversal and inversion symmetry breaking are known to be necessary conditions that do not guarantee the onset of a diode effect. In turn, casting the analysis in terms of mirror symmetry breaking is more stringent. Although it is only possible to demonstrate that junctions that do not break the relevant mirror symmetry cannot show JDE, all the cases studied do indeed show JDE.
@@ IV Acknowledgments
We acknowledge S. Heun for valuable discussions and for carefully reading the manuscript. L.C. acknowledges the Fondazione Cariplo under the grant 2023-2594. M.C. and F.G. acknowledge the EU’s Horizon 2020 Research and Innovation Framework Programme under Grants No. 964398 (SUPERGATE) and the PNRR MUR project PE0000023-NQSTI for partial financial support. F.G. acknowledges partial financial support under Grant No. 101057977 (SPECTRUM). E.S. acknowledges the project HelicS, project number DFM.AD002.206. We acknowledge NextGenerationEU PRIN project 2022A8CJP3 (GAMESQUAD) for partial financial support.